OpenStudy (guy5000):

Burning one gallon of gasoline releases 20 pounds of CO2 into the air. Harry’s compact car can travel 25 miles on one gallon of gas. If Harry drives 15,000 miles in a year, his vehicle will add pounds of CO2 to the atmosphere.

OpenStudy (mathmate):

@guy5000 Your post does not look like a question. In case you just missed out the question mark at the last sentence, and you'd like to calculate the number of pounds (#) of CO2 per year by Harry's car: This kind of seemingly confusing problem is much easier if we use units in the calculations. Given: 20 lbs/gallon 25 miles /gallon 15000 miles /year We want lbs/year, so 20 lbs/gallon \(\div\) 25 miles/gal \(\times\) 15000 miles /year =\(20\times \frac{1}{25} \times 15000\) lbs/gallon * gallon/mile * miles/year =\(20\times \frac{1}{25} \times 15000\) lbs/year (after cancelling units) I'll let you work it out with or without a calculator.